OpenStudy (anonymous):

What is the value of f(4) for the function f(x) = 3x – 5?

OpenStudy (anonymous):

my answer is 2

hartnn (hartnn):

to get f(4) just replace 'x' in f(x) by 4

hartnn (hartnn):

so what is 3*4-5 =.... ?

OpenStudy (nurali):

f(4) = 3(4) – 5 f(4) = 12-5 f(4)=_________

OpenStudy (anonymous):

7

hartnn (hartnn):

yes

OpenStudy (nurali):

Correct.

OpenStudy (anonymous):

what year are you?

OpenStudy (anonymous):

thank you. love you guys

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

This is an extra course